OpenStudy (anonymous):

Find the slope and the y-intercept of the equation 4x + 5y = -9.

OpenStudy (anonymous):

slope intercept equation of a straight line, \[y=mx+c\] \[m=slope\]\[c=y intercept\] isolate y and make it's coefficient equal to unity

OpenStudy (anonymous):

The slope in this equation is 4, because in slope intercept form {y=mx+b} M= slope and b= y intercept :)

OpenStudy (anonymous):

@E>Nattie12345 Please don't give direct answers, rather guide the students to their answers

OpenStudy (anonymous):

I explained the answer...that's not the same :)

OpenStudy (anonymous):

i have a comprehension disability

OpenStudy (anonymous):

@E>Nattie12345 Your answer is incorrect I believe anyway

OpenStudy (anonymous):

@joshfoster200 I'll help you with the first step josh but you must do the next \[4x+5y=-9\] y on one side \[4x+5y-4x=-9-4x\]\[5y=-4x-9\] Next step is to make the number right next to y equal to 1 of the form \[y=mx+c\]

OpenStudy (anonymous):

The slope is not 4! Change to slope-intercept form before identifying slope and y-intercept.